# One missing piece in Vision and Language: A Survey on Comics Understanding

### v1

@@ 6.1.3 Action Detection
Unique to Action Detection is its multimodal input nature, where each panel in the sequence contributes to a holistic understanding of the action. Unlike single image analysis, this task requires the model to interpret a series of images, each adding context and continuity to the narrative. Despite being popular for Videos [93], this task has not yet been explored in comics apart from image-cloze task [8] where the model is challenged to select the correct next panel from two options. This task, while elementary, underscores the potential for more sophisticated Action Detection in comics, a domain where sequential imagery plays a pivotal role.
6.1.4 Page Stream Segmentation
Definition: Page Stream Segmentation (PSS) in comics entails segmenting a comic book into different stories or sections. In its simplistic version, the task can be seen as categorizing each page with a tag among cover, regular page, ads, or last page. In its more advanced form, it includes extracting salient information from the stories such as title, authors, drawer, price, characters, and synopsis.
The concept of PSS, while not new in document analysis, has seen various methodologies employed over time. Traditional approaches, such as those utilizing Support Vector Machines (SVM) for single page classification [94], have evolved with the advent of Convolutional Neural Networks (CNNs) [95], and more recently, attention-based multimodal fusion techniques [96]. These methods have been pivotal in processing and understanding page layouts and structures in various document formats.
Despite its longstanding presence in document analysis, PSS remains largely uncharted in the comics domain. Comics present a unique challenge for PSS due to their complex structure; a single volume can interweave multiple stories, interspersed with advertisements and narrative text. This intricacy demands a nuanced approach to segment and understand the diverse content within comic books.
A potential advancement in PSS for comics involves not just segmenting the pages but also tagging these segments with metadata like story titles, authors, and synopses. This represents a promising field of study, extending traditional image classification into a more intricate, story-centric context.
6.2 Augmentation
Augmentation tasks, in the context of comics, involve image manipulation techniques aimed at enhancing or altering the visual presentation. The primary tasks identified under this category are Super-Resolution (SR), Style-Transfer (ST), Vectorization, and Depth Estimation. Each of these tasks revolves around processing a single image input to produce an enhanced or stylistically altered output.
6.2.1 Super-Resolution
@@
Recent advancements in manga super-resolution (SR) focus on preserving visual integrity during enhancement. Yao et al. [97] employ deep learning to classify and tailor SR models for manga screentone, ensuring semantic preservation. Dai et al. [98] introduce the Structured Fusion Attention Network (SFAN), which optimizes feature extraction and reconstruction quality through attention modules, achieving superior performance on datasets like Manga109. These methods enhance image quality while maintaining the artistic nuances critical to manga, effectively balancing technical efficiency with artistic integrity.
6.2.2 Style-Transfer
@@
Photo-to-Comic Transformation: Photo-to-comic has mainly involved mangas and involves replacing colors and textures with halftone patterns. Traditionally, it often consists of manually selecting and adjusting the colors. Methods have evolved from basic pattern generation to sophisticated techniques capturing structural and tonal similarities. Classic approaches relied on algorithmic solutions and were limited in their ability to handle multiple screentone. Recent advancements by Zhang et al. [99] utilize deep learning for semantic region segmentation in cartoon images, followed by screentone application. However, these methods are primarily effective for color-based inputs, with limited applicability to line-based images.
@@
Image-to-Image Translation: Topal et al. [100] explored style transfer as a transfer for the detection task. Specifically, they used style transfer networks like CycleGAN and CartoonGAN to train a detection network, which was then fine-tuned on specific comic datasets for character and face detection. This approach, to use style transfer techniques to enhance performances in other tasks, is not new [68, 101].
Figure 5: Colorization vs. Screening
Colorization: Colorization in comics, specifically in manga, represents the task of transforming black-and-white images into full-color pages. It is a complex process that requires a balance of artistic understanding and technical precision.
Furusawa et al. [102] introduced a comprehensive approach to the colorization of manga. Their method allows for user interaction, where color dots can be added to guide the colorization process. This approach not only automates the colorization but also gives users the flexibility to influence the final output, ensuring that the automatic colorization aligns with the intended aesthetic vision. In a different approach, Hensman et al. [103] presented a model based on Conditional Generative Adversarial Networks (cGANs). Their model, specifically tuned for each reference image, colorizes manga panels to mimic colorization styles. This method implies training a new GAN for each reference image, making it versatile for different styles but requiring significant computational resources for each new style adaptation. Subsequently, in [104] authors investigated screentone synthesis, a critical aspect of manga art that involves translating line drawings into screentone-patterned images. They utilized Manga109 for comparison with standard image-to-image translation methods like pix2pix [105]. Their approach, although similar in architecture to U-Net used in previous studies [106], stands out by explicitly considering screentone labels, thus preserving the integrity of screentone patterns in the final output. Moving towards a more advanced solution, Golyadkin et al. [107] introduced a two-stage architecture for manga page colorization. The first stage involves drafting the coloring using the Pixel2Style2Pixel architecture [108], followed by a conditional GAN in the second stage to refine the output. This method elegantly fuses color drafts with color hints from users, generating high-quality colorizations. They demonstrated improvements both qualitatively and quantitatively, indicating a significant advancement in manga colorization technology. Building upon previous methods, Jiramahapokee et al. [109] further refined the colorization process. They leveraged a multi-encoder Variational Autoencoder (VAE) for improved color region consistency and combined it with the shading strengths of other tools like manga-colorization-v2. Their approach also incorporated CIELAB [110] interpolation to enhance color saturation and authenticity, leading to outputs that are not only aesthetically pleasing but also consistent and detailed in their shading quality.
@@
Screening: Screening, or adding screentone to black-and-white manga drawings, has been a focus of several studies. A notable contribution by Wu et al. [111] introduced a two-stage architecture for this process. The first stage involves generating grayscale shading using a six-layer Swin-transformer [112], which then serves as input for the second stage, a reference-based screentone generation module. This algorithmic process extracts screentone patterns from a reference manga and applies them to line drawings, considering the shading nuances provided by the first stage.
@@ 6.2.3 Vectorization
Structural Lines Extraction: This process, structural lines extraction, focuses on removing screentones to leave only the defining lines of figures and shapes. It can be seen as the reverse operation of the screening, which corresponds to adding the screentone to a structural line image. In this task, Li et al. [106] developed a CNN-based method using a U-Net architecture with residual blocks to effectively extract structural lines from manga pages. This method is adaptable for various applications, including manga retargeting and colorization, and can handle diverse patterns and textures, resulting in clean skeleton sketches.
Manga Vectorization: Various works [113, 114, 115, 116] have proposed different approaches for this task. In particular, Yao et al. [117] developed a method involving adaptive binarization and screentone detection, refining borders, estimating lighting, and compensating for missing strokes, leading to high-quality, resolution-independent rendering. A recent innovative approach, MARVEL by Su et al. [118], employs Reinforcement Learning to train an agent in selecting drawing primitives that recreate raster manga images in vector format. While effective, this approach works with squared manga patches and may result in visual gaps in certain renderings. Moreover, this method is not convenient in shape rendering, as it uses primitives and cannot manipulate them. MARVEL represents an interesting exploration into the use of machine learning for vectorization, highlighting both the potential and the complexities of this task.
6.2.4 Depth Estimation
@@
Up to now, this necessity arose particularly when adapting comics for various digital media, such as smartphones, tablets, and e-readers, where the original layout may not fit the display dimensions optimally. The primary methodologies for reconfiguring comic panels involve: (i) Adjusting panel proportions and rescaling them to fit the new medium without compromising the narrative or visual integrity. (ii) Strategically cropping panels while ensuring that key objects within each panel are preserved and remain visible in the adapted format. Both techniques rely heavily on depth estimation to identify and prioritize the most significant elements in each panel.
One significant challenge in this domain is the lack of ground truth data for comics. To address this, Bhattacharjee et al. [7] adopted an innovative approach using Image-to-Image style-transfer. They transformed comic images to resemble natural images, upon which depth estimation techniques, involving Laplacian edges and feature-based GANs, were applied to predict depth effectively. Building on this, a subsequent study [119], based on their earlier work [120], introduced a Swin Transformer for multitasking, including both depth estimation and object segmentation. This multi-decoder-head architecture, while effective, presents challenges in scalability due to the complexity of accommodating multiple tasks.
Depth Estimation plays a pivotal role in adapting comics to the varying formats of digital consumption. By ensuring that key visual elements are retained and appropriately displayed, these methodologies not only enhance the reader’s experience but also preserve the artistic intent of the original work.
6.3 Satellite Tasks
Some of the tasks that fit under the Augmentation umbrella have not found a space in our taxonomy, e.g. the task of de-warping. The task involves correcting distortions typically found in scanned comic pages. Addressing the de-warping challenge, a recent study by Garai et al. [121] proposed a novel mathematical model to describe the warping process. This model estimates distortion factors based on the boundaries of panels in a comic document image, aiming to rectify these warps effectively. While their approach yielded promising qualitative results, particularly in the image areas, challenges persist in handling text distortion, where some distortions proved unrecoverable.
7 Layer 2: Grounding, Analysis, and Segmentation
@@
7.1 Grounding
Grounding in comics involves identifying and classifying specific elements within a comic’s panel, such as text, characters, and objects. Under this task name umbrella, many individual tasks belong such as Detection, Character Re-identification, and sentence Grouding.
7.1.1 Detection
@@
Panels. Initial studies in digital Manga, such as those by Ponsard et al. [122] and Arai et al. [123], emphasized automated segmentation and panel detection, underscoring the importance of maintaining reading order. Arai et al. [124] further developed these ideas with a focus on real-time detection and extraction techniques, which evolved into more sophisticated SVG-based approaches as explored in [125]. [126] contributed advanced methods for panel extraction using region growing techniques and mathematical morphology, combined with speech balloon detection.
@@
Text. Jomaa et al. [127] employed a tracking algorithm for panel extraction, while speech balloons were identified using Robert’s edge detection operator and classified with thresholding. Liu et al. [128] focused on balloon extraction based on text character locations. Pal et al. [129] performed text extraction using SIFT local features, Space Pyramid Matching (SPM), and SVM on eBDtheque and Bangla comics images. In [130] authors contributed to text extraction in Manga109 combining region proposal generation with SVM classifiers. Chu et al. [131] and Qin et al. [34] modified existing architectures like Faster R-CNN for detecting text and faces in comics. In particular, [131] proposed additional aspect ratio regions, while [34] built two comics datasets, namely JC2463 (Manga black and white style) and AEC912 (colored American comics style) to test face detection evaluation.
Face. Interestingly, in [132] authors benchmarked Fast R-CNN, Faster R-CNN and SSD in panel, character, face, and balloon detection. They used the Manga109 dataset, discovering that Fast R-CNN works better in panels when there are clear boundaries, Faster R-CNN when boundaries are not well defined, and SSD needs already cut panels as it is limited for full-page complex mangas.
@@
So far, many methods have used classical detection architectures and applied them to the comic domain to extract elements such as panels, characters, text, etc. A new direction is drawn by [133] challenged the conventional four-edge polygon representation of comic panels, suggesting segmentation-based approaches using U-Net for more accurate extraction. Here, authors performed extraction through segmentation, employing a U-Net that classifies pixels into background, panels, and borders.
@@
Onomatopoeia. Recently, the focus has shifted to onomatopoeia detection, a relatively unexplored area in comic analysis. The release of the COO dataset by Baek et al. [4] has paved the way for new research in this domain. Authors proposed a modified version of M4C [134] designed to link truncated onomatopoeias together. Following this, Yang et al. [92] proposed text augmentation techniques for onomatopoeias using EfficientNet.
Sharma et al. [135] furthered this research by introducing their dataset for panel and character detection, employing a Faster R-CNN model. However, they did not provide comparative analysis with other models.
@@
Benchmarks. Lately, Dutta et al. [136, 69] introduced the BCBId dataset, benchmarking existing datasets (eBDtheque, Manga109, and DCM) and employing YOLO architectures for panel and character detection. They presented comparisons with existing methods, showing enhanced performance in detecting panels and characters across multiple datasets. However, it is unclear if these results are truly comparable. This is due to the fact that in previous (and contemporary) works, authors don’t provide neither information about the settings used (in terms of train-val-split), nor about hyperparameters and model settings to replicate the weights. Only recently, Vivoli et al. [2] fairly compares models across various datasets and styles, unifying annotations for multiple detection classes and providing model weights, code, and data to replicate.
7.1.2 Character Re-Identification

### v2

@@ 6.1.3. Action Detection
Unique to Action Detection is its multimodal input nature, where each panel in the sequence contributes to a holistic understanding of the action. Unlike single image analysis, this task requires the model to interpret a series of images, each adding context and continuity to the narrative. Despite being popular for Videos (Zhu et al., 2020), this task has not yet been explored in comics apart from image-cloze task (Iyyer et al., 2016) where the model is challenged to select the correct next panel from two options. In the context of clozure, many works have proposed valid solutions, starting from LSTM-based original solution of (Iyyer et al., 2016), the BERT-style multimodal and multigranularity embedding of (Soykan et al., 2024a), and the Multimodal T5-based approach of (Vivoli et al., 2024d). However, while elementary, this task underscores the potential for more sophisticated Panel sequence analysis (and Action Detection) in comics, a domain where sequential imagery plays a pivotal role. Recently, (Ikuta et al., 2024) propose panel-level benchmarks about text and character recognition, as well as visual-cloze tasks in manga-style comics benchmarking recent open-source MLLMs models.
6.1.4. Page Stream Segmentation
Definition: Page Stream Segmentation (PSS) in comics involves dividing a comic book into different sections or stories. In its simplest form, this entails categorizing pages with tags like cover, regular page, ads, or last page. In a more advanced form, PSS extracts key information such as story titles, authors, illustrators, prices, characters, and synopses.
Though PSS is not a new concept in document analysis, it has seen various methods evolve over time. Traditional approaches like Support Vector Machines (SVM) for single-page classification (Gordo et al., 2013) have given way to Convolutional Neural Networks (CNNs) (Wiedemann and Heyer, 2017) and more recent attention-based multimodal fusion techniques (Demirtacs et al., 2022), which have significantly improved the understanding of page layouts and structures in documents. However, PSS remains largely unexplored in the comics domain, where the task is complicated by the unique structure of comics. A single volume may contain multiple interwoven stories, ads, and narrative text, creating a more complex segmentation challenge. This complexity requires more sophisticated methods for segmenting and understanding the diverse content of comic books. An emerging area of interest in PSS for comics is not just segmenting the pages but also tagging each segment with metadata such as story titles, authors, and synopses.
Figure 6. Illustration of Augmentation tasks.
This expands traditional image classification into a more detailed, story-oriented approach, offering a promising field of study.
6.2. Augmentation
Augmentation tasks, in the context of comics, involve image manipulation techniques aimed at enhancing or altering the visual presentation (Fig. 6). The primary tasks identified under this category are Super-Resolution (SR), Style-Transfer (ST), Vectorization, and Depth Estimation. Each of these tasks revolves around processing a single image input to produce an enhanced or stylistically altered output.
6.2.1. Super-Resolution
@@
Recent advancements in manga super-resolution (SR) focus on preserving visual integrity during enhancement. Yao et al. (Yao et al., 2023) employ deep learning to classify and tailor SR models for manga screentone, ensuring semantic preservation. Dai et al. (Dai et al., 2022) introduce the Structured Fusion Attention Network (SFAN), which optimizes feature extraction and reconstruction quality through attention modules, achieving superior performance on datasets like Manga109. These methods enhance image quality while maintaining the artistic nuances critical to manga, effectively balancing technical efficiency with artistic integrity.
6.2.2. Style-Transfer
@@
Photo-to-Comic Transformation: Photo-to-comic has mainly involved mangas and involves replacing colors and textures with halftone patterns. Traditionally, it often consists of manually selecting and adjusting the colors. Methods have evolved from basic pattern generation to sophisticated techniques capturing structural and tonal similarities. Classic approaches relied on algorithmic solutions and were limited in their ability to handle multiple screentone. Recent advancements by Zhang et al. (Zhang et al., 2021) utilize deep learning for semantic region segmentation in cartoon images, followed by screentone application. However, these methods are primarily effective for color-based inputs, with limited applicability to line-based images.
@@
Image-to-Image Translation: Topal et al. (Topal et al., 2022) explored style transfer as a transfer for the detection task. Specifically, they used style transfer networks like CycleGAN and CartoonGAN to train a detection network, which was then fine-tuned on specific comic datasets for character and face detection. This approach -using style transfer techniques to enhance performances in other tasks- has also been explored in different applications (Inoue et al., 2018; Deng et al., 2020). In particular, the multi-task learning method for dense predictions in comic panels has been used as a medium for comic panel reconfiguration (Bhattacharjee et al., 2023) and depth estimation (Bhattacharjee et al., 2021).
Colorization: Colorization in comics, particularly manga, involves transforming black-and-white images into full-color pages, requiring both artistic precision and technical expertise. Furusawa et al. (Furusawa et al., 2017) introduced a method that allows users to add color dots to guide an automated process, giving users the flexibility to influence the final output while ensuring the system aligns with their vision. Hensman et al. (Hensman and Aizawa, 2017) used Conditional GANs (cGANs) to colorize manga panels by tailoring the model to each reference image, allowing versatile style adaptations but requiring a new GAN for every image, increasing computational demands. Tsubota et al. (Tsubota et al., 2019) explored
Figure 7. Colorization vs. Screening
screentone synthesis, essential for manga, by comparing image translation techniques like pix2pix (Isola et al., 2016) on the Manga109 dataset. Their approach, using a U-Net architecture (Li et al., 2017), stands out by explicitly considering screentone labels, preserving the screentone patterns integral to manga art. Golyadkin et al. (Golyadkin and Makarov, 2023) introduced a two-stage method for manga page colorization: Pixel2Style2Pixel (Richardson et al., 2020) for color drafts, followed by a conditional GAN to refine the output. This method elegantly combines user-provided color hints and draft colorizations, yielding high-quality results that surpass previous methods. Jiramahapokee et al. (Jiramahapokee, 2023) further refined the process using a multi-encoder Variational Autoencoder (VAE) to ensure color consistency and integrated CIELAB interpolation (Zeyen et al., 2018) to enhance color saturation and authenticity. Their approach produced detailed and visually consistent colorizations, particularly in shading quality.
@@
Screening: Screening, or adding screentone to black-and-white manga drawings, has been a focus of several studies. A notable contribution by Wu et al. (Wu et al., 2023) introduced a two-stage architecture for this process. The first stage involves generating grayscale shading using a six-layer Swin-transformer (Liu et al., 2021a), which then serves as input for the second stage, a reference-based screentone generation module. This algorithmic process extracts screentone patterns from a reference manga and applies them to line drawings, considering the shading nuances provided by the first stage.
@@ 6.2.3. Vectorization
Structural Lines Extraction: This process, structural lines extraction, focuses on removing screentones to leave only the defining lines of figures and shapes. It can be seen as the reverse operation of the screening, which corresponds to adding the screentone to a structural line image. In this task, Li et al. (Li et al., 2017) developed a CNN-based method using a U-Net architecture with residual blocks to effectively extract structural lines from manga pages. This method is adaptable for various applications, including manga retargeting and colorization, and can handle diverse patterns and textures, resulting in clean skeleton sketches.
Manga Vectorization: Various works (Kopf and Lischinski, 2011; Selinger, 2003; Ma et al., 2022) have proposed different approaches for this task. In particular, Yao et al. (Yao et al., 2017) developed a method involving adaptive binarization and screentone detection, refining borders, estimating lighting, and compensating for missing strokes, leading to high-quality, resolution-independent rendering. A recent innovative approach, MARVEL by Su et al. (Su et al., 2024), employs Reinforcement Learning to train an agent in selecting drawing primitives that recreate raster manga images in vector format. While effective, this approach works with squared manga patches and may result in visual gaps in certain renderings. Moreover, this method is not convenient in shape rendering, as it uses primitives and cannot manipulate them. MARVEL represents an interesting exploration into the use of machine learning for vectorization, highlighting both the potential and the complexities of this task.
6.2.4. Depth Estimation
@@
To date, this need has arisen primarily when adapting comics for digital media like smartphones, tablets, and e-readers, where the original layout may not fit the screen optimally. Key methods for reconfiguring panels include: (i) Adjusting panel proportions and rescaling them to fit the new medium while maintaining narrative and visual integrity, and (ii) Cropping panels to preserve essential elements. Both approaches rely on depth estimation to identify and prioritize key components. A major challenge in this area is the lack of ground truth data for comics. To address this, Bhattacharjee et al. (Bhattacharjee et al., 2021) employed Image-to-Image style-transfer, transforming comic images to resemble natural ones, enabling depth estimation using Laplacian edges and feature-based GANs. Building on this, a follow-up study (Bhattacharjee et al., 2023), based on their earlier work (Bhattacharjee et al., 2022), introduced a Swin Transformer for multitasking, including depth estimation and object segmentation. However, the multi-decoder-head architecture poses scalability challenges due to task complexity. Depth estimation is crucial for adapting comics to various digital formats, ensuring key visual elements are retained while preserving the original artistic intent.
6.3. Satellite Tasks
Some of the tasks that fit under the Augmentation umbrella have not found a space in our taxonomy, e.g. the task of de-warping. The task involves correcting distortions typically found in scanned comic pages. Addressing the de-warping challenge, a recent study by Garai et al. (Garai et al., 2022) proposed a novel mathematical model to describe the warping process. This model estimates distortion factors based on the boundaries of panels in a comic document image, aiming to rectify these warps effectively. While their approach yielded promising qualitative results, particularly in the image areas, challenges persist in handling text distortion, where some distortions proved unrecoverable.
7. Layer 2: Grounding, Analysis, and Segmentation
@@
7.1. Grounding
Figure 8. Illustration of Grounding tasks.
Grounding in comics involves identifying and classifying specific elements within a comic’s panel, such as text, characters, and objects (Fig. 8). Under this task name umbrella, many individual tasks belong such as Detection, Character Re-identification, and sentence Grouding.
7.1.1. Detection
@@
Panels. Initial studies in digital Manga, such as those by Ponsard et al. (Ponsard, 2009) and Arai et al. (Arai and Tolle, 2010), emphasized automated segmentation and panel detection, underscoring the importance of maintaining reading order. Arai et al. (Arai and Tolle, 2011) further developed these ideas with a focus on real-time detection and extraction techniques, which evolved into more sophisticated SVG-based approaches as explored in (Su et al., 2011). (Ho et al., 2012) contributed advanced methods for panel extraction using region growing techniques and mathematical morphology, combined with speech balloon detection.
@@
Text. Jomaa et al. (Jomaa et al., 2015) employed a tracking algorithm for panel extraction, while speech balloons were identified using Robert’s edge detection operator and classified with thresholding. Liu et al. (Liu et al., 2015b) focused on balloon extraction based on text character locations. Pal et al. (Pal et al., 2016) performed text extraction using SIFT local features, Space Pyramid Matching (SPM), and SVM on eBDtheque and Bangla comics images. In (Aramaki et al., 2016) authors contributed to text extraction in Manga109 combining region proposal generation with SVM classifiers. Chu et al. (Chu and Yu, 2018) and Qin et al. (Qin et al., 2017) modified existing architectures like Faster R-CNN for detecting text and faces in comics. In particular, (Chu and Yu, 2018) proposed additional aspect ratio regions, while (Qin et al., 2017) built two comics datasets, namely JC2463 (Manga black and white style) and AEC912 (colored American comics style) to test face detection evaluation.
@@
Face. Interestingly, in (Yanagisawa et al., 2018) authors benchmarked Fast R-CNN, Faster R-CNN and SSD in panel, character, face, and balloon detection. They used the Manga109 dataset, discovering that Fast R-CNN works better in panels when there are clear boundaries, Faster R-CNN when boundaries are not well defined, and SSD needs already cut panels as it is limited for full-page complex mangas. So far, many methods have used classical detection architectures and applied them to the comic domain to extract elements such as panels, characters, text, etc. A new direction is drawn by (Nguyen et al., 2019b) challenged the conventional four-edge polygon representation of comic panels, suggesting segmentation-based approaches using U-Net for more accurate extraction. Here, authors performed extraction through segmentation, employing a U-Net that classifies pixels into background, panels, and borders.
@@
Onomatopoeia. Recently, the focus has shifted to onomatopoeia detection, a relatively unexplored area in comic analysis. The release of the COO dataset by Baek et al. (Baek et al., 2022) has paved the way for new research in this domain. Authors proposed a modified version of M4C (Hu et al., 2019) designed to link truncated onomatopoeias together. Following this, Yang et al. (Yang and Chu, 2023) proposed text augmentation techniques for onomatopoeias using EfficientNet. Sharma et al. (Sharma and Kukreja, 2023) furthered this research by introducing their dataset for panel and character detection, employing a Faster R-CNN model. However, they did not provide comparative analysis with other models.
@@
Benchmarks. Lately, Dutta et al. (Dutta and Biswas, 2019; Dutta et al., 2022) introduced the BCBId dataset, benchmarking existing datasets (eBDtheque, Manga109, and DCM) and employing YOLO architectures for panel and character detection. They presented comparisons with existing methods, showing enhanced performance in detecting panels and characters across multiple datasets. However, it is unclear if these results are truly comparable. This is due to the fact that in previous (and contemporary) works, authors don’t provide neither information about the settings used (in terms of train-val-split), nor about hyperparameters and model settings to replicate the weights. Only recently, Vivoli et al. (Vivoli et al., 2024c) fairly compare models across various datasets and styles, unifying annotations for multiple detection classes and providing model weights, code, and data to replicate.
7.1.2. Character Re-Identification